For saving energy, bicycling and walking are far more efficient means of transportation than is travel by automobile. For example, when riding at 12.0 mi/h, a cyclist uses food energy at a rate of about 420 kcal/h above what he would use if merely sitting still. (In exercise physiology, power is often measured in kcal/h rather than in watts. Here 1 kcal = 1 nutritionist's Calorie = 4186 J.) Walking at 2.80 mi/h requires about 210 kcal/h. It is interesting to compare these values with the energy consumption required for travel by car. Gasoline yields about 1.30 x10^8 J/gal.
(a) Find the fuel economy in equivalent miles per gallon for a person walking.
(b) Find the fuel economy in equivalent miles per gallon for a person bicycling.
1
Expert's answer
2019-11-06T09:10:02-0500

A) For a person walking:

find the amount of energy spent walking

210*4186=879060 J,

find the part of the amount of energy from the amount of energy corresponding to 1 gallon of fuel

879060/130000000=0.00672

find the number of miles corresponding to a gallon of fuel, taking into account the speed of the pedestrian

2.8/0.006762~414 m/g


B) for a person bicycling:

find the amount of energy spent riding a bike


420*4186=1758120 J

find the part of the amount of energy from the amount of energy corresponding to 1 gallon of fuel 1758120/130000000=0.013525

find the number of miles corresponding to a gallon of fuel, taking into account the speed of the cyclist


12/0.013524~887 m/g

Answer: a) 414 m/g, b)887 m/g